If the difference between the square of a positive integer and the integer before it is 33, what is the integer?
Answer by ankor@dixie-net.com(22740)

If the difference between the square of a positive integer and the integer before it, is 31, what is the integer?
:
x^2 - (x-1) = 31
x^2 - x + 1` = 31
x^2 - x + 1 - 31 = 0
x^2 - x - 30 = 0
Factors to
(x-6)(x+5) = 0
The positive solution
x = 6 is the integer
:
;
See if that works 6^2 - 5 = 31